# Algebraically Skew Embedding of Curves

### v1

@@ 4.2 The factorization of $\widetilde{D}_{1}\to D_{1}$
$c_{1}(\mathcal{T}_{\widetilde{D}_{1}})=\pi_{D}^{*}(4\zeta_{1}+4\zeta_{2}+7e)-{j_{D}}_{*}(2).$ (33)
4.3 The class $\left\llbracket D_{1}^{\dagger}\right\rrbracket$
@@
$[\widetilde{\Delta}_{\Gamma}]=$ $(\widetilde{\pi}_{\Delta}^{*}([\hat{\gamma}(X)]))\zeta_{\mathcal{H}}^{2}+\left(\widetilde{\pi}_{\Delta}^{*}(c_{1}(\mathsf{Hom}(\mathcal{S}_{2}/\mathcal{S}_{1},\mathcal{V}/\mathcal{S}_{2}))[\hat{\gamma}(X)])+\hat{\gamma}_{*}(K)\right)\zeta_{\mathcal{H}}$ (34)
@@
$c_{1}(\mathsf{Hom}(\mathcal{S}_{2}/\mathcal{S}_{1},\mathcal{V}/\mathcal{S}_{2}))=$ $c_{1}\left((\mathcal{S}_{2}/\mathcal{S}_{1})^{*}\otimes(\mathcal{V}/\mathcal{S}_{2})\right)$ (35)
@@
$[\hat{\gamma}(X)]=$ $de^{3}(\zeta_{\mathcal{Q}}-e)^{3}+((5e)(de^{3})+(2g-2)e^{4})(\zeta_{\mathcal{Q}}-e)^{2}$ (36)
@@
$\left\llbracket i_{D}^{\dagger*}c_{1}(\mathcal{T}_{\widetilde{D}_{1}})\right\rrbracket=(15d+20g-20)P,$ (37)
@@
$\left\llbracket i^{\dagger*}c_{1}(\mathcal{T}_{B})\right\rrbracket=(10d^{\vee}+10g-10)P.$ (38)
@@
4.4 The class $\left\llbracket\Gamma^{\dagger}\right\rrbracket$ and the product $\left\llbracket D_{1}^{\dagger}\right\rrbracket\left\llbracket\Gamma^{\dagger}\right\rrbracket$
@@
$j^{\dagger}_{*}[E^{\dagger}\cap\Gamma^{\dagger}]=$ $j^{\dagger}_{*}\left((\zeta^{\dagger})^{10}+\pi_{E}^{\dagger*}c_{1}\left(\mathcal{N}_{\widetilde{\Delta}_{\Gamma},B}/\mathcal{N}_{\widetilde{\Delta}_{\Gamma},\widetilde{\Gamma}}\right)(\zeta^{\dagger})^{9}\right)$ (39)
@@
$\left\llbracket D_{1}^{\dagger}\right\rrbracket\left\llbracket\Gamma^{\dagger}\right\rrbracket=$ $(d^{\vee})^{2}P-\pi^{\dagger*}j_{*}\left\llbracket 5d^{\vee}\overline{\sigma}_{33}\zeta^{5}-15d^{\vee}\overline{\sigma}_{33}\zeta^{5}+(15d^{\vee}+6g-6)\overline{\sigma}_{33}\zeta^{5}\right\rrbracket$ (40)
@@
4.5 The class $\left\llbracket\varepsilon_{\Delta^{\dagger}_{\Gamma}}\right\rrbracket$
@@
$\left\llbracket c_{1}(\mathcal{T}_{\Delta^{\dagger}_{\Gamma}})\right\rrbracket=(2-2g)P$ (41)
@@
$\left\llbracket\iota_{\Delta^{\dagger}_{\Gamma}\Gamma^{\dagger}}^{*}[\Gamma^{\dagger}]\right\rrbracket=\left\llbracket[\Delta_{X}](-K\otimes 1-1\otimes K)\right\rrbracket=4-4g.$ (42)
@@
$\iota_{\Delta^{\dagger}_{\Gamma}}^{*}\left\llbracket c_{1}(\mathcal{T}_{B^{\dagger}})\right\rrbracket=(10d^{\vee}+30g-30)P,$ (43)
@@
$\iota_{\Delta^{\dagger}_{\Gamma}D_{1}^{\dagger}}^{*}\left\llbracket c_{1}(\mathcal{T}_{D_{1}^{\dagger}})\right\rrbracket=(15d+36g-36)P.$ (49)
@@
$\left\llbracket\varepsilon_{\Delta^{\dagger}_{\Gamma}}\right\rrbracket=$ $\left\llbracket\iota_{\Delta^{\dagger}_{\Gamma}}^{*}c_{1}(\mathcal{T}_{B^{\dagger}})\right\rrbracket-\left\llbracket\iota_{\Delta^{\dagger}_{\Gamma}\Gamma^{\dagger}}^{*}c_{1}(\mathcal{T}_{\Gamma^{\dagger}})\right\rrbracket-\left\llbracket\iota_{\Delta^{\dagger}_{\Gamma}D^{\dagger}_{1}}^{*}c_{1}(\mathcal{T}_{D^{\dagger}_{1}})\right\rrbracket+\left\llbracket c_{1}(\mathcal{T}_{\Delta^{\dagger}_{\Gamma}})\right\rrbracket$ (50)
@@
4.6 The skew curves in $\mathbb{P}^{4}$
@@
We have verified this number with Macaulay 2. Note that this number, curiously, coincides with the cardinality of a Steiner complex a genus $5$ curve, see [Dol12, Proposition 5.4.7].
@@
By the nondegeneracy of $X$, we have $d\geq 4$. Together with the quadratic equation above, we see that $6\leq d^{\vee}\leq 16$ and $d\leq 5$. Furthermore, since both $d$ and $d^{\vee}$ are integers, the remaining options are $(d,d^{\vee})=(4,6),(5,10),(5,12),(4,16)$. In other words, the only possible cases are
$g$ 00 $1$ $2$ $5$
$d$ $4$ $5$ $5$ $4$.
However, since these are nondegenerate smooth curves in $\mathbb{P}^{4}$, the cases $(5,2),(4,5)$ are in fact impossible, see [GH94, p.253]. Therefore, the remaining possible cases are the rational normal curve with $d=4$ and $g=0$, and the elliptic normal curves with $d=5$ and $g=1$.
@@
Note that by taking the affine part of an elliptic normal curves in $\mathbb{P}^{4}=\mathbb{CP}^{4}$, we obtain a skew embedding of a $5$-punctured torus into $\mathbb{R}^{8}$, which gives us a nontrivial real totally skew surface in an Euclidean space of dimension $<9$.

### v4

@@ 4.4. The factorization of $\widetilde{D}_{1}\to D_{1}$
(33) $c_{1}(\mathcal{T}_{\widetilde{D}_{1}})=\pi_{D}^{*}(4\zeta_{1}+4\zeta_{2}+7e)-{j_{D}}_{*}(2).$
4.5. The class $\left\llbracket D_{1}^{\dagger}\right\rrbracket$
@@
(34) $[\widetilde{\Delta}_{\Gamma}]=$ $(\widetilde{\pi}_{\Delta}^{*}([\hat{\gamma}(X)]))\zeta_{\mathcal{H}}^{2}+\left(\widetilde{\pi}_{\Delta}^{*}(c_{1}(\mathsf{Hom}(\mathcal{S}_{2}/\mathcal{S}_{1},\mathcal{V}/\mathcal{S}_{2}))[\hat{\gamma}(X)])+\hat{\gamma}_{*}(K)\right)\zeta_{\mathcal{H}}$
@@
(35) $c_{1}(\mathsf{Hom}(\mathcal{S}_{2}/\mathcal{S}_{1},\mathcal{V}/\mathcal{S}_{2}))=$ $c_{1}\left((\mathcal{S}_{2}/\mathcal{S}_{1})^{*}\otimes(\mathcal{V}/\mathcal{S}_{2})\right)$
@@
(36) $[\hat{\gamma}(X)]=$ $de^{3}(\zeta_{\mathcal{Q}}-e)^{3}+((5e)(de^{3})+(2g-2)e^{4})(\zeta_{\mathcal{Q}}-e)^{2}$
@@
(37) $\left\llbracket i_{D}^{\dagger*}c_{1}(\mathcal{T}_{\widetilde{D}_{1}})\right\rrbracket=(15d+20g-20)P,$
@@
(38) $\left\llbracket i^{\dagger*}c_{1}(\mathcal{T}_{B})\right\rrbracket=(10d^{\vee}+10g-10)P.$
@@
4.6. The class $\left\llbracket\Gamma^{\dagger}\right\rrbracket$ and the product $\left\llbracket D_{1}^{\dagger}\right\rrbracket\left\llbracket\Gamma^{\dagger}\right\rrbracket$
@@
(39) $j^{\dagger}_{*}[E^{\dagger}\cap\Gamma^{\dagger}]=$ $j^{\dagger}_{*}\left((\zeta^{\dagger})^{10}+\pi_{E}^{\dagger*}c_{1}\left(\mathcal{N}_{\widetilde{\Delta}_{\Gamma},B}/\mathcal{N}_{\widetilde{\Delta}_{\Gamma},\widetilde{\Gamma}}\right)(\zeta^{\dagger})^{9}\right)$
@@
(40) $\left\llbracket D_{1}^{\dagger}\right\rrbracket\left\llbracket\Gamma^{\dagger}\right\rrbracket=$ $(d^{\vee})^{2}P-\pi^{\dagger*}j_{*}\left\llbracket 5d^{\vee}\overline{\sigma}_{33}\zeta^{5}-15d^{\vee}\overline{\sigma}_{33}\zeta^{5}+(15d^{\vee}+6g-6)\overline{\sigma}_{33}\zeta^{5}\right\rrbracket$
@@
4.7. The class $\left\llbracket\varepsilon_{\Delta^{\dagger}_{\Gamma}}\right\rrbracket$
@@
(41) $\left\llbracket c_{1}(\mathcal{T}_{\Delta^{\dagger}_{\Gamma}})\right\rrbracket=(2-2g)P$
@@
(42) $\left\llbracket\iota_{\Delta^{\dagger}_{\Gamma}\Gamma^{\dagger}}^{*}[\Gamma^{\dagger}]\right\rrbracket=\left\llbracket[\Delta_{X}](-K\otimes 1-1\otimes K)\right\rrbracket=4-4g.$
@@
(43) $\iota_{\Delta^{\dagger}_{\Gamma}}^{*}\left\llbracket c_{1}(\mathcal{T}_{B^{\dagger}})\right\rrbracket=(10d^{\vee}+30g-30)P,$
@@
(49) $\iota_{\Delta^{\dagger}_{\Gamma}D_{1}^{\dagger}}^{*}\left\llbracket c_{1}(\mathcal{T}_{D_{1}^{\dagger}})\right\rrbracket=(15d+36g-36)P.$
@@
(50) $\left\llbracket\varepsilon_{\Delta^{\dagger}_{\Gamma}}\right\rrbracket=$ $\left\llbracket\iota_{\Delta^{\dagger}_{\Gamma}}^{*}c_{1}(\mathcal{T}_{B^{\dagger}})\right\rrbracket-\left\llbracket\iota_{\Delta^{\dagger}_{\Gamma}\Gamma^{\dagger}}^{*}c_{1}(\mathcal{T}_{\Gamma^{\dagger}})\right\rrbracket-\left\llbracket\iota_{\Delta^{\dagger}_{\Gamma}D^{\dagger}_{1}}^{*}c_{1}(\mathcal{T}_{D^{\dagger}_{1}})\right\rrbracket+\left\llbracket c_{1}(\mathcal{T}_{\Delta^{\dagger}_{\Gamma}})\right\rrbracket$
@@
4.8. The skew curves in $\mathbb{P}^{4}$
@@
We have verified this number with Macaulay 2. Note that this number, curiously, coincides with the cardinality of a Steiner complex of a genus $5$ curve, see [Dol12, Proposition 5.4.7]. The relation between theta characteristics and tangents of canonical curves has been observed in [BC23] as well as [Cil24], but the specific relation to the Steiner complex remains unknown to the author.
@@
By the nondegeneracy of $X$, we have $d\geq 4$. Together with the quadratic equation above, we see that $6\leq d^{\vee}\leq 16$ and $d\leq 5$. Furthermore, since both $d$ and $d^{\vee}$ are integers, the remaining options are $(g,d)=(0,4),(1,5),(2,5),(5,4)$. However, since these are nondegenerate smooth curves in $\mathbb{P}^{4}$, the cases $(5,2),(4,5)$ are in fact not possible, see [GH94, p.253]. Therefore, the remaining possible cases are the rational normal curve with $d=4$ and $g=0$, and the elliptic normal curves with $d=5$ and $g=1$.
@@
Note that by taking the affine part of an elliptic normal curve $X\subseteq\mathbb{P}^{4}=\mathbb{CP}^{4}$ with the infinity chosen to be a hyperplane intersecting $X$ at one point with multiplicity $5$, we obtain a skew embedding of a punctured torus into $\mathbb{R}^{8}$, which gives us a nontrivial real totally skew surface in an Euclidean space of dimension $<9$.